OpenStudy (anonymous):

Medal A mother gives birth to triplets. The individual weights of the babies at birth are 4 lb 8 oz, 4 lb 6 oz, and 3 lb 8 oz. Find the total birth weight of the triplets. ? lb ? oz

OpenStudy (mathmate):

1 lb = ? oz

OpenStudy (mathmate):

@harz360 !

OpenStudy (anonymous):

sorry lb oz

OpenStudy (mathmate):

I suppose you mean 16 oz in 1 lb. So 4 lb 6oz equals 4*16+6=70 oz. Do the same for the others, add all three together, and divide by 3 to get the average. Convert back to lbs and oz, and that's your answer.

OpenStudy (anonymous):

i still dont understand

OpenStudy (mathmate):

Example: Three cakes weight 2 lb 8oz, 1 lb 3oz, 2 lb 5 oz. Step 1: Convert all to the smaller unit, oz. 2 lb 8oz = 2*16+8=40 oz 1 lb 3oz = 1*16+3 = 19 oz 2 lb 5 oz = 2*16+5 = 37 oz 2. Calculate sum (in smaller unit) sum = 40+19+37 = 96 oz 3. Calculate average average = sum/number of cakes = 96 oz / 3 cakes = 32 oz/cake 4. Convert to appropriate units 1 lb=16 oz so 32 oz = 32/16 = 2 lb 0oz 5. Give answer in words The average weight of each cake is 2 lb 0 oz each.